# Geometry-aware RL for Manipulation of Varying Shapes and Deformable Objects

### v2

@@ Abstract
Manipulating objects with varying geometries and deformable objects is a major challenge in robotics. Tasks such as insertion with different objects or cloth hanging require precise control and effective modelling of complex dynamics. In this work, we frame this problem through the lens of a heterogeneous graph that comprises smaller sub-graphs, such as actuators and objects, accompanied by different edge types describing their interactions. This graph representation serves as a unified structure for both rigid and deformable objects tasks, and can be extended further to tasks comprising multiple actuators. To evaluate this setup, we present a novel and challenging reinforcement learning benchmark, including rigid insertion of diverse objects, as well as rope and cloth manipulation with multiple end-effectors. These tasks present a large search space, as both the initial and target configurations are uniformly sampled in 3D space. To address this issue, we propose a novel graph-based policy model, dubbed Heterogeneous Equivariant Policy (HEPi), utilizing $SE(3)$ equivariant message passing networks as the main backbone to exploit the geometric symmetry. In addition, by modeling explicit heterogeneity, HEPi can outperform Transformer-based and non-heterogeneous equivariant policies in terms of average returns, sample efficiency, and generalization to unseen objects.
@@ 2 Background
In this work, we focus on geometric graphs ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\mathcal{X}})$, where each node $v$ is associated with coordinates ${\bm{x}}_{v}\in{\mathcal{X}}=\mathbb{R}^{3}$ and steerable geometric features $\tilde{\mathbf{f}}_{v}$. A feature is considered steerable if it transforms consistently under the action of a group $g\in G$ through a group representation $\rho$. For instance, a vector ${\bm{v}}\in\mathbb{R}^{3}$ under a rotation ${\bm{R}}\in SO(3)$ transforms as ${\bm{v}}^{\prime}={\bm{R}}{\bm{v}}$. In this case, ${\bm{v}}$ is a finite-dimensional vector, and its transformation is described by an invertible matrix representation with determinant $1$, $\rho(g)={\bm{R}}$.
@@
A Markov Decision Process (MDP) is defined by the tuple $(S,A,P,R,\gamma)$, where $S$ is the set of states, $A$ the actions, $P(s^{\prime}|s,a)$ the transition probability, $R(s,a)$ the reward function, and $\gamma\in[0,1]$ the discount factor. The goal is to find a policy $\pi:S\to A$ that maximizes the expected discounted reward $E[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$.
In MDPs with symmetries, both the transition function $P(s^{\prime}|s,a)$ and policy function $\pi(a|s)$ are equivariant under group transformations $g\in G$ via left-regular representation $L_{g}$ and $K_{g}$ for state and action, respectively:
@@
while the reward function $R(s,a)$ remains invariant: $R(L_{g}[s],K_{g}^{s}[a])=R(s,a)$. This allows leveraging symmetries to reduce the complexity of learning, potentially improving sample efficiency and generalization, as it results in a group-structured MDP homomorphism (Van der Pol et al., 2020).
@@ 3.1 Equivariant MPN Backbone
An Equivariant Message Passing Network (EMPN) can be constructed (Brandstetter et al., 2022; Duval et al., 2023; Bekkers et al., 2024) by enforcing equivariance in the functions $\phi$ and $\psi$ in Equation 1, ensuring that their inputs and outputs are steerable and transform consistently under the group $G$, as described in Definition 2.1. Constructing such functions for high-dimensional steerable features is challenging and typically requires spherical harmonics embeddings (Duval et al., 2023), where matrix-vector multiplications are carried out using Clebsch-Gordan tensor products, followed by steerable activation functions (Brandstetter et al., 2022; Bekkers et al., 2024). While these operations guarantee equivariance, they also introduce high computational complexity, making them impractical for reinforcement learning settings. To mitigate this issue, Bekkers et al. (2024) introduced the PONITA framework, an efficient equivariant message-passing approach. In this work, we adopt PONITA as our EMPN backbone and, for consistency, refer to it as EMPN throughout the rest of the paper.
Consider the message function in Equation 1, where $\psi(\mathbf{f}_{v}^{(k)},\mathbf{f}_{u}^{(k)},\mathbf{e}_{uv})=k({\bm{x}}_{u}-{\bm{x}}_{v})\mathbf{f}_{u}$ is defined as a linear function with steerable feature $\mathbf{f}_{u}$, and summation is used as the aggregation function, $\bigoplus$=$\sum$ = ∑. Here, $k$ is a convolution kernel, and $({\bm{x}}_{u}-{\bm{x}}_{v})$ represents the relative position between nodes $u$ and $v$. This leads to the following convolutional message-passing update rule:
@@
On a regular grid, e.g., an image, each relative position $({\bm{x}}_{u}-{\bm{x}}_{v})$ has a corresponding weight ${\bm{W}}_{u,v}$ and is stored in one single matrix ${\bm{W}}$. However, this does not apply to non-uniform grids, e.g. point clouds, $k({\bm{x}}_{u},{\bm{x}}_{v})$ in this case can be parameterized by a neural network. To ensure equivariance, the kernel function $k({\bm{x}}_{u},{\bm{x}}_{v})$ must transform consistently under the group action $g\in G$.
To this end, features in EMPN are embedded in both position and orientation spaces using a uniform grid on $\mathcal{S}^{2}$. It “lifts” the domain from ${\mathcal{X}}=\mathbb{R}^{3}$ to $\mathcal{X}^{\uparrow}=\mathbb{R}^{3}\times\mathcal{S}^{2}$. Hence, the kernel function in this position-orientation space is defined as
$k([({\bm{p}},{\bm{o}}),({\bm{p}}^{\prime},{\bm{o}}^{\prime})])=K^{(3)}\,k^{(2)}({\bm{o}}^{\top}{\bm{o}}^{\prime})\,k^{(1)}({\bm{o}}^{\top}({\bm{p}}^{\prime}-{\bm{p}}),\|{\bm{o}}\perp({\bm{p}}^{\prime}-{\bm{p}})\|),$
where $k^{(1)}$ handles spatial interactions based on the relative position $({\bm{p}}^{\prime}-{\bm{p}})$ and the perpendicular component $\|{\bm{o}}\perp({\bm{p}}^{\prime}-{\bm{p}})\|$, $k^{(2)}$ manages orientation-based interactions via dot products ${\bm{o}}^{\top}{\bm{o}}^{\prime}$, and $K^{(3)}$ performs channel-wise mixing across features.
By factorizing the kernel into spatial, orientation, and channel components, the operations remain efficient without requiring specialized network structures. Although this discretization sacrifices exact equivariance, it drastically reduces computational costs. Furthermore, while the EMPN was originally designed for homogeneous graphs, its structure can be adapted to tasks involving heterogeneous graphs, as we explore in the next section.
@@ 3.2 Heterogeneous Equivariant Policy
$\theta_{k+1}=\arg\max_{\theta}\mathbb{E}_{(s,a)\sim\pi_{\theta_{k}}}\left[\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{k}}(a|s)}A^{\pi_{\theta_{k}}}(s,a)\right]\quad\text{s.t.}\quad\bar{D}_{\text{KL}}(\pi_{\theta}||\pi_{\theta_{k}})\leq\delta,$
where $A^{\pi_{\theta_{k}}}(s,a)$ is the advantage function. Here, $\bar{D}_{\text{KL}}(\pi_{\theta}||\pi_{\theta_{k}})$ is the KL-divergence between the new policy $\pi_{\theta}$ and the old policy $\pi_{\theta_{k}}$, constrained by $\delta$ to ensure stable updates and prevent overly large policy changes. PPO approximates this trust region by clipping the importance sampling ratio to limit updates. This, however, requires careful hyperparameters turning to make it work stably, as pointed out by Andrychowicz et al. (2021). We will show later in the Results Section 4.2, this is also applied to graph-based policy. On the other hand, Trust Region Projection Layers (TRPL) (Otto et al., 2021) adopt a more principled approach. TRPL projects policy parameters onto trust region boundaries using a differentiable convex optimization, ensuring stability by projecting both the mean and variance of the Gaussian policy to satisfy trust region constraints.
@@ Acknowledgments
We also thank Niklas Freymuth, Onur Celik, and Aleksandar Taranovic for insightful discussions during the initial and rebuttal phases of the paper. Additionally, we appreciate Reviewer 6nhr for suggesting the Rigid-Pushing task.

### v6

@@ Abstract
Manipulating objects with varying geometries and deformable objects is a major challenge in robotics. Tasks such as insertion with different objects or cloth hanging require precise control and effective modelling of complex dynamics. In this work, we frame this problem through the lens of a heterogeneous graph that comprises smaller sub-graphs, such as actuators and objects, accompanied by different edge types describing their interactions. This graph representation serves as a unified structure for both rigid and deformable objects tasks, and can be extended further to tasks comprising multiple actuators. To evaluate this setup, we present a novel and challenging reinforcement learning benchmark, including rigid insertion of diverse objects, as well as rope and cloth manipulation with multiple end-effectors. These tasks present a large search space, as both the initial and target configurations are uniformly sampled in 3D space. To address this issue, we propose a novel graph-based policy model, dubbed Heterogeneous Equivariant Policy (HEPi), utilizing $SE(3)$ equivariant message passing networks as the main backbone to exploit the geometric symmetry. In addition, by modeling explicit heterogeneity, HEPi can outperform Transformer-based and non-heterogeneous equivariant policies in terms of average returns, sample efficiency, and generalization to unseen objects. Our project page is available here.
@@ 2 Background
In this work, we focus on geometric graphs ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\mathcal{X}})$, where each node $v$ is associated with coordinates ${\bm{x}}_{v}\in{\mathcal{X}}=\mathbb{R}^{3}$ and steerable geometric features $\mathbf{f}_{v}$. A feature is considered steerable if it transforms consistently under the action of a group $g\in G$ through a group representation $\rho$. For instance, a vector ${\bm{v}}\in\mathbb{R}^{3}$ under a rotation ${\bm{R}}\in SO(3)$ transforms as ${\bm{v}}^{\prime}={\bm{R}}{\bm{v}}$. In this case, ${\bm{v}}$ is a finite-dimensional vector, and its transformation is described by an invertible matrix representation with determinant $1$, $\rho(g)={\bm{R}}$.
@@
A Markov Decision Process (MDP) is defined by the tuple $(S,A,P,R,\gamma)$, where $S$ is the set of states, $A$ the actions, $P(s^{\prime}|s,a)$ the transition probability, $R(s,a)$ the reward function, and $\gamma\in[0,1]$ the discount factor. The goal is to find a policy $\pi:S\to A$ that maximizes the expected discounted reward $\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$.
In MDPs with symmetries, both the transition distribution $P(s^{\prime}|s,a)$ and policy distribution $\pi(a|s)$ are invariant under group transformations $g\in G$ via left-regular representation $L_{g}$ and $K_{g}$ for state and action, respectively, resulting in the following conditions:
@@
and similarly for the reward function: $R(L_{g}[s],K_{g}^{s}[a])=R(s,a)$. This allows leveraging symmetries to reduce the complexity of learning, potentially improving sample efficiency and generalization, as it results in a group-structured MDP homomorphism (Van der Pol et al., 2020).
@@ 3.1 Equivariant MPN Backbone
An Equivariant Message Passing Network (EMPN) can be constructed (Brandstetter et al., 2022; Duval et al., 2023; Bekkers et al., 2024) by enforcing equivariance in the functions $\phi$ and $\psi$ in Equation 1, ensuring that their inputs and outputs are steerable and transform consistently under the group $G$, as described in Definition 2.1. Constructing such functions for high-dimensional steerable features is challenging and typically requires spherical harmonics embeddings (Duval et al., 2023), where matrix-vector multiplications are carried out using Clebsch-Gordan tensor products, followed by steerable activation functions (Brandstetter et al., 2022; Bekkers et al., 2024). While these operations guarantee equivariance, they also introduce high computational complexity, making them impractical for reinforcement learning settings. To mitigate this issue, Bekkers et al. (2024) introduced the PONITA framework, an efficient equivariant message-passing approach. We use it as our EMPN backbone and refer to it as EMPN throughout the rest of the paper for consistency.
@@
[…]
Furthermore, to improve computational efficiency, the kernel function is factorized as:
$k_{\theta}([({\bm{p}}_{u},{\bm{o}}_{u}),({\bm{p}}_{v},{\bm{o}}_{v})])=K_{\theta}^{(3)}\,k_{\theta}^{(2)}({\bm{o}}_{v}^{\top}{\bm{o}}_{u})\,k_{\theta}^{(1)}({\bm{o}}_{v}^{\top}({\bm{p}}_{u}-{\bm{p}}_{v}),\|{\bm{o}}_{v}\perp({\bm{p}}_{u}-{\bm{p}}_{v})\|),$
where $k^{(1)}$ handles spatial interactions based on the relative position $({\bm{p}}_{u}-{\bm{p}}_{v})$ and the perpendicular component $\|{\bm{o}}_{v}\perp({\bm{p}}_{u}-{\bm{p}}_{v})\|$, $k^{(2)}$ manages orientation-based interactions via dot products ${\bm{o}}_{v}^{\top}{\bm{o}}_{u}$, and $K^{(3)}$ performs channel-wise mixing across features. This formulation preserves the universal approximation property of equivariant functions while being significantly more computationally efficient and not requiring specialized network structures. Furthermore, ${\bm{o}}$ can be sampled on a uniform grid over $\mathcal{S}^{2}$, making EMPN only approximately equivariant. However, in practice, it achieves strong performance even with a limited number of samples, as discussed in Appendix D.
@@ 3.2 Heterogeneous Equivariant Policy
$\theta_{k+1}=\arg\max_{\theta}\mathbb{E}_{(s,a)\sim\pi_{\theta_{k}}}\left[\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{k}}(a|s)}A^{\pi_{\theta_{k}}}(s,a)\right]\quad\text{s.t.}\quad D_{\text{KL}}(\pi_{\theta}||\pi_{\theta_{k}})\leq\delta,$
where $A^{\pi_{\theta_{k}}}(s,a)$ is the advantage function. Here, $D_{\text{KL}}(\pi_{\theta}||\pi_{\theta_{k}})$ is the KL-divergence between the new policy $\pi_{\theta}$ and the old policy $\pi_{\theta_{k}}$, constrained by $\delta$ to ensure stable updates and prevent overly large policy changes. PPO approximates this trust region by clipping the importance sampling ratio to limit updates. This, however, requires careful hyperparameters turning to make it work stably, as pointed out by Andrychowicz et al. (2021). We will show later in the Results Section 4.2, this is also applied to graph-based policy. On the other hand, Trust Region Projection Layers (TRPL) (Otto et al., 2021) adopt a more principled approach. TRPL projects policy parameters onto trust region boundaries using a differentiable convex optimization, ensuring stability by projecting both the mean and variance of the Gaussian policy to satisfy trust region constraints.
@@ Acknowledgments
We thank Niklas Freymuth, Philipp Dahlinger, Onur Celik, Aleksandar Taranovic, and Mayank Mittal for their insightful discussions during both the initial and rebuttal phases of this paper, as well as for their technical support. We also appreciate Reviewer 6nhr for suggesting the Rigid-Pushing task.